OpenStudy (anonymous):

-9(t-6)-(t+3)=8 The solution is t=?

OpenStudy (ryan123345):

\[-9t+54-t-3=8\]

OpenStudy (ryan123345):

\[-10t+51=8\]

OpenStudy (ryan123345):

\[-10t=-43\]

OpenStudy (ryan123345):

t=43/10

OpenStudy (anonymous):

the answer is t=43/10?